OpenStudy (anonymous):

The sides of a rectangle are in the ratio 5:7 and the perimeter of the rectangle is 96cm What is the area of the rectangle?

OpenStudy (anonymous):

2x+2y=96 y=(5/7)x 2x+(10/7)x=96 (24/7)x=96 x=96*7/24 x=28 y=20 Area =560